OpenStudy (minecraftqueen):

need help ASAP will medal and fan What is the slope-intercept form of the function that contains the point (8, 12) and has a slope of −2? y =-2x +6 is this right

OpenStudy (anonymous):

umm i think -4

OpenStudy (anonymous):

12=-2(8)+b add 2 to both sides 14=(8)+b subtract 8 from both sides 6=b

OpenStudy (minecraftqueen):

-_- wait what @mjnnnnnnnnnn

OpenStudy (anonymous):

its -4

OpenStudy (anonymous):

i think im not positive